A wire of resistance R is bent in form of a square with side a what is magnetic induction at centre due to current in wire 
 

Grade:12

1 Answers

Pooja
4 years ago
We calculate flux in unit time which is magnetic induction
 
Flux = BA
 
For loop length is a so A=a2
Resistance of full length is R
As it shaped into square the resistance of each arm would be R/4  so current in each arm would be I/4 the original current
 
B= 4× field due to one wire at centre
   = 4 ×  uI/4 ÷ 4 π a/2 × {2÷(2)1/2}
   = I/(2)1/2
Induction at centre =I/ (2)1/2a × a2
        = Ia/(2)1/2